Circular No. 8469

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005L IN MCG +07-33-5 T. Boles, Coddenham, Suffolk, England, reports the discovery of an

apparent supernova (mag 18.2) on unfiltered CCD images taken on Jan. 15.141 and 19.073 UT with a 0.35-m reflector. The new object is located at α = 15^{h}53^{m}25^{s}.92, δ = +39^{o}23′08′′.1, which is ≈ 5′′.1 east and 2′′.7 south of the center of MCG +07-33-5. Nothing is present at this position on Boles’ images from 2004 June 5 and Sept. 5 (limiting mag 19.5) or on Digitized Sky Survey plates from 1991 (limiting red mag 21.0) and 1988 (limiting blue mag 20.5).

COMETS C/2004 X10, X11, Y2, Y3, Y4 (SOHO) Further to IAUC 8466, following are the initial observations of addi-

tional comets found on SOHO website images (GG = G. Gallina; HO = H. Otterstedt); all are Kreutz sungrazers except for C/2004 Y4, which has no known group membership and reached mag 5–6 on Dec. 26 (but showed no tail).

Comet  2004 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2004 X10 Dec. 13.292 17^{h}27.^{m}9 −25^{o}05′ C2 KB 2005-A41
C/2004 X11  13.883 17 30.1 −25 03 C2 GG 2005-A41
C/2004 Y2  18.975 17 53.7 −25 18 C2 HS 2005-A41
C/2004 Y3  24.279 18 22.1 −28 37 C3 TS 2005-A41
C/2004 Y4  24.679 18 14.6 −17 32 C3 HO 2005-A41

COMET C/2005 B1 = 2004 FS_{101} (CHRISTENSEN) An object reported as asteroidal by LINEAR (2004 Mar. 23) and by

Spacewatch (2004 Mar. 18 and 26), and published as 2004 FS_{101} (MPS 104028, 110712; initial orbit on MPEC 2004-O45 and MPO 66351) has been identified with C/2005 B1. The following revised orbital elements are taken from MPEC 2005-B25:

Epoch = 2006 Mar. 6.0 TT

T = 2006 Feb. 23.4751 TT  ω = 103.^{o}1841
e = 1.000709  Ω = 195.5570 }2000.0
q = 3.204423 AU  i = 92.5535
2005 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green